Find the distinct interval of length 1 containing a root or solutin of f(x) = x³ - 3x + 5 using IVT

1
Expert's answer
2022-03-18T12:33:05-0400

Solution


The given function is


"f(x) = x\u00b3 - 3x + 5"


"f(-4) = (-4)\u00b3 - 3(-4) + 5 = -47<0"


"f(0) = (0)\u00b3 - 3(0) + 5 = 5>0"


Since "f(0)>0" and "f(-4)<0" , therefore, the root of the given function lies between the interval "[-4, 0]"


The length of this interval "[-4, 0]" is "=(0)-(-4)=4" .


We narrow this interval, such that we get the length of the interval equal to 1.


"f(-3) = (-3)\u00b3 - 3(-3) + 5 = -13<0"


"f(-1) = (-1)\u00b3 - 3(-1) + 5 =7 >0"


Again we see that "f(-1)>0" and "f(-3)<0" , therefore, the root of the given function lies between the interval "[-3, -1]"


The length of this interval "[-3, -1]" is "=(-1)-(-3)=2" .


Now consider,


"f(-2) = (-2)\u00b3 - 3(-2) + 5 =3 >0"


Still we can see that t "f(-2)>0" and "f(-3)<0" , therefore, the root of the given function lies between the interval "[-3, -2]"


And the length of the interval is "=(-2)-(-3)=1".



Hence the required interval of length 1, where the root of the given function lies is [−3,−2]



The diagram below is the plot of the given function "f(x) = x\u00b3 - 3x + 5" , which entirely ensures that the root lies within the interval [−3,−2]





The root is "-2.279", which is within the interval "[-3,-2]"